# Sharp Monocular View Synthesis in Less Than a Second

### v1

@@ 1 Introduction
We conduct a thorough experimental evaluation on multiple datasets that were not used during training, using powerful perceptual metrics such as LPIPS (Zhang et al., 2018) and DISTS (Ding et al., 2022) to assess image fidelity. SHARP improves image fidelity by substantial factors versus prior feed-forward methods. In comparison to diffusion-based systems, SHARP delivers higher fidelity while reducing synthesis time by two to three orders of magnitude. Compared to the strongest prior method (Ren et al., 2025), SHARP reduces LPIPS by 25–34% and DISTS by 21–43% across the test datasets (in the zero-shot regime), while accelerating synthesis by three orders of magnitude and producing a 3D representation that supports high-resolution rendering of nearby views at 100 frames per second.
In summary, our contributions are as follows
@@
Using our insights, we demonstrate that state-of-the-art high-resolution view synthesis is feasible in a purely regression based framework.
@@ 2 Related Work
A number of works develop feed-forward prediction models for view synthesis from a small number of nearby views. IBRNet (Wang et al., 2021) generalized image-based rendering across scenes using learned features and ray transformers. MVSNeRF (Chen et al., 2021) reconstructed neural radiance fields from a few input images via cost volume processing. LaRa (Chen et al., 2024) regressed an object-level radiance field from sparse input images. GS-LRM (Zhang et al., 2024) leveraged a transformer to predict a 3D Gaussian representation from posed sparse images. Our work focuses on view synthesis from a single image.
View synthesis from a single image. The introduction of larger datasets (Zhou et al., 2018; Tongue et al., 2024) enabled the transition from scene-specific multi-view optimization to learning-based pipelines that can infer a plausible 3D representation from a single image. Zhou et al. (2016) synthesized novel views from a single image through appearance flow. Subsequent work developed variants of depth-based warping (Wiles et al., 2020; Jampani et al., 2021) and multiplane images (MPI) (Zhou et al., 2018; Tucker & Snavely, 2020). AdaMPI (Han et al., 2022) adapted multiplane images to diverse scene layouts through plane depth adjustment and depth-aware color prediction, trained using a warp-back strategy on single-view image collections. Khan et al. (2023) proposed Tiled Multiplane Images (TMPI), which splits an MPI into many small tiled regions with fewer depth planes per tile, reducing computational overhead while maintaining quality. Several recent methods have drawn inspiration from the success of transformers in modeling long-range dependencies, leveraging large transformer-based encoder-decoder architectures to infer scene structure and appearance implicitly for novel views (Hong et al., 2023; Jin et al., 2025).
@@
Diffusion models have emerged as powerful tools for novel view synthesis with sparse input, offering high-quality results through iterative denoising processes (Po et al., 2023). Watson et al. (2022) developed an early application to view synthesis. Zero-1-to-3 (Liu et al., 2023) demonstrated zero-shot view synthesis by fine-tuning diffusion models on 3D object datasets. iNVS (Kant et al., 2023) repurposed diffusion inpainters for view synthesis by combining monocular depth estimation with inpainting. NerfDiff (Gu et al., 2023) distilled a 3D-aware diffusion model into NeRF by synthesizing virtual views to improve rendering under occlusion. These early attempts focused on object-level view synthesis.
At the scene level, ViewCrafter (Yu et al., 2024), ZeroNVS (Sargent et al., 2024), CAT3D (Gao et al., 2024), SplatDiff (Zhang et al., 2025), Stable Virtual Camera (Zhou et al., 2025), Bolt3D (Szymanowicz et al., 2025b), Wonderland (Liang et al., 2025), WonderWorld (Yu et al., 2025), See3D (Ma et al., 2025) and Gen3C (Ren et al., 2025) all applied diffusion models to view synthesis from sparse image sets or a single image. The diffusion-based approach supports impressive image quality from faraway viewpoints, leveraging diffusion priors to synthesize plausible appearance even for views that have no overlap with the input. On the other hand, image quality from nearby views (corresponding to natural head motion or posture shifts) can be noticeably less sharp and photorealistic than the input, while the synthesis time can sometimes stretch into minutes. (Although Bolt3D (Szymanowicz et al., 2025b) makes impressive progress on the latter front.) In contrast, we aim for real-time rendering of maximally photorealistic high-resolution images from nearby views, supporting a headbox that allows for natural posture shifts while maintaining photographic quality. Our approach generates a high-resolution 3D representation that provides such experiences from single-image input in less than a second on a single GPU, supporting conversion of pre-existing photographs to photorealistic 3D during interactive browsing of a photo collection.
@@ 3.1 Overview
The input image $\mathbf{I}\in\mathbb{R}^{3\times H\times W}$ is fed into a pretrained Depth Pro image encoder $\varphi_{\text{enc}}$ to produce 4 intermediate features maps $(\mathbf{f}_{i})_{i\in 1,\dots,4}=\varphi_{\text{enc}}(\mathbf{I})$. As in Depth Pro (Bochkovskii et al., 2025), we resize the input image so that $H=W=1536$. We then feed the intermediate feature maps into the Depth Pro decoder $\varphi_{\text{dec}}$ to produce a monocular depth map $\mathbf{\hat{D}}=\varphi_{\text{dec}}((\mathbf{f}_{i})_{i\in 1,\dots,4})$. Similar to Flynn et al. (2019), we duplicate the last convolutional layer of the decoder to produce a two-channel depth map $\mathbf{\hat{D}}\in\mathbb{R}^{2\times H\times W}$.
@@
While the initial Gaussians provide a reasonable starting point, they require substantial refinement to achieve high-fidelity rendering. The Gaussian decoder $\varphi_{\text{gauss}}$ takes as input the feature maps $(\mathbf{f}_{i})_{i\in 1,\dots,4}$ and the input image $\mathbf{I}$, and outputs refinements $\Delta\mathbf{G}\in\mathbb{R}^{K\times 2\times H^{\prime}\times W^{\prime}}$} for all Gaussian attributes:
$\Delta\mathbf{G}=\varphi_{\text{gauss}}((\mathbf{f}_{i})_{i\in 1,\dots,4},\mathbf{I}).$ (3.1)
@@
$\mathbf{G}_{\text{attr}}=\gamma_{\text{attr}}\bigg(\gamma_{\text{attr}}^{-1}(\mathbf{G}_{0,\text{attr}})+\eta_{\text{attr}}\Delta\mathbf{G}_{\text{attr}}\bigg),$ (3.2)
@@ 3.2 Network Architecture
Feature encoder. We base our feature encoder on the Depth Pro backbone (Bochkovskii et al., 2025). This encoder processes the input image $\mathbf{I}\in\mathbb{R}^{C\times H\times W}$} and produces four feature maps $(\mathbf{f}_{i})_{i\in 1,\dots,4}$ at different resolutions. The Depth Pro backbone consists of two Vision Transformers (ViTs) (Dosovitskiy et al., 2021), one applied to a downscaled version of the input image and one applied to various image patches. The low-resolution image encoder and patch encoder each has 326M parameters. During training, we unfreeze the low-resolution image encoder to allow adaptation to the view synthesis task, while keeping the patch encoder and normalization layers frozen to preserve the pretrained feature extraction capabilities.
Depth decoder. Our depth decoder is based on the Dense Prediction Transformer (DPT) (Ranftl et al., 2021). We modify the original DPT decoder by duplicating the final convolutional layer to output two depth channels instead of one. Our encoder thus takes the feature maps $(\mathbf{f}_{i})_{i\in 1,\dots,4}$ from the encoder and produces a two-layer depth map $\mathbf{\hat{D}}\in\mathbb{R}^{2\times H\times W}$. The first layer represents the primary visible surfaces, while the second layer may represent occluded regions and view-dependent effects. The depth decoder consists of multiple convolutional blocks with approximately 20M parameters. This module is fully unfrozen during training to optimize depth prediction for view synthesis.
Gaussian decoder. The Gaussian decoder predicts refinements for all Gaussian attributes. It has the same DPT architecture as the depth decoder but we replace the last upsampling block with a custom prediction head. The decoder takes as input the feature maps $(\mathbf{f}_{i})_{i\in 1,\dots,4}$, the input image $\mathbf{I}$, and the predicted depth maps $\mathbf{\hat{D}}$. It outputs a tensor $\Delta\mathbf{G}\in\mathbb{R}^{K\times 2\times H^{\prime}\times W^{\prime}}$ that contains deltas for all Gaussian attributes: position (3 channels), scale (3 channels), rotation (4 channels), color (3 channels), and opacity (1 channel). This decoder has approximately 7.8M parameters and is trained from scratch. The high dimensionality of the output (approximately 16.5M values) enables fine-grained control over the Gaussian representation.
@@ 3.4 Training Objectives
where $\Omega$ denotes the set of all pixels $p$. We further use a perceptual loss (Johnson et al., 2016; Gatys et al., 2016; Suvorov et al., 2021) on novel views to encourage plausible inpainting:
@@
$\mathcal{L}_{\text{delta}}=\mathbb{E}_{i\sim\mathcal{I}}\left[\max\{\left|\Delta\mathbf{G}_{x}(i)\right|-\delta,0\}+\max\{\left|\Delta\mathbf{G}_{y}(i)\right|-\delta,0\}\right]$ (3.9)
@@
$\mathcal{L}_{\text{scale}}=\mathbb{E}_{p\sim\Omega}\left[|\mathbf{S}(p)|\right]\quad\text{ and }\quad\mathcal{L}_{\nabla\text{scale}}=\sum_{k=1}^{6}\mathbb{E}_{p\sim\Omega_{\downarrow k}}\left[|\nabla\mathbf{S}_{\downarrow k}(p)|\right].$ (3.11)
@@ 4 Experiments
We employ LPIPS (Zhang et al., 2018) and DISTS (Ding et al., 2022) to quantitatively assess the quality of novel view synthesis. We focus primarily on these perceptual metrics, since older pointwise metrics such as PSNR and SSIM can be overly sensitive to small translations, where even a 1% shift can lead to catastrophic drops in scores despite visually similar results. (See the supplement for an illustration. We also list PSNR and SSIM numbers in the supplement for completeness.) Since we are interested in sharp high-resolution view synthesis, we evaluate all methods on the full-resolution ground truth. If a method generates results at lower resolution, we resize the output image to the input resolution before evaluation. If a method crops the input and generates cropped results (Ren et al., 2025; Yu et al., 2024; Jin et al., 2025; Zhou et al., 2025), we evaluate against correspondingly cropped ground-truth images.
@@
We compare SHARP to the following state-of-the-art methods: Flash3D (Szymanowicz et al., 2025a), which is based on 3D Gaussians; TMPI (Khan et al., 2023), which uses multi-plane images; LVSM (Jin et al., 2025), which is based on image-to-image regression; and Stable Virtual Camera (SVC) (Zhou et al., 2025), ViewCrafter (Yu et al., 2024), and Gen3C (Ren et al., 2025), which employ diffusion models.
@@ 5 Conclusion
One clear opportunity for future work is to extend the methodology to support photorealistic synthesis of faraway views without compromising the fidelity of nearby views or the benefits of fast interactive synthesis. This may call for judicious integration of diffusion models (Po et al., 2023), possibly with the aid of distillation for reducing synthesis latency (Yin et al., 2024). With diffusion models, a unified view synthesis routine for single view, multi-view, and video input (Ren et al., 2025; Ma et al., 2025) may emerge as a versatile generalization. Another interesting avenue is a principled treatment of view-dependent and volumetric effects (Verbin et al., 2024).
@@ A.1 Attribute Specific Activation
Activation functions $\gamma$ and their correspondent scale factors $\eta$ in Eq. 3.2 are specified below:
@@ A.3 View Frustum Masking
We implement a view frustum masking technique to address ambiguity in view synthesis – regions occluded in the original view have multiple plausible reconstructions. By using depth information to determine which regions in the new view correspond to points visible in the original view, we apply supervision only where ground truth is reliable.
@@
A.4 Perceptual Loss
Here we detail the challenges and our solutions in incorporating the perceptual loss. We employ the perceptual loss aimed at improving inpainting (Suvorov et al., 2021). Similar to its application in the image domain, we initially applied the loss only in the occluded image patches of the novel view; however, through experiments, we observed that applying the loss to the entire rendered image resulted in more plausible details and fewer artifacts in general, even in the non-occluded foreground regions, as seen in Figure 9.
@@ B.1 Synthetic data
To further enhance scene diversity and complexity, the framework supports random placement of various object types, including thin structures, transparent materials, and reflective surfaces, across a wide range of spatial configurations. It also offers fine-grained control over camera parameters such as position, orientation, and focal length, as well as detailed illumination settings. Lighting setups include both physically-based direct light sources, with variations in direction, intensity, and color temperature. We also use high-dynamic-range (HDR) environment maps, which are sampled from a curated collection off high-resolution HDRIs. This combination enables realistic global illumination effects under diverse and physically plausible lighting conditions.
For each scene, we identify one object of interest and position a ring containing 10 virtual cameras around it at varying distances and angles. The cameras are arranged in concentric circles such that the cameras no more than 60cm apart, simulating a multi-view capture setup. This allows the dataset to capture the same object or scene element from diverse perspectives. All images are rendered using the V-Ray physically based rendering engine, ensuring photorealistic lighting and material interactions. The final dataset consists of approximately 700K unique rendered scene instances, each with 11 rendered views, totaling around 8M images at $1536\times 1536$ or $2048\times 2048$ resolutions.

### v2

# arXiv:2512.10685v2[cs.CV] 27 Feb 2026

@@ 1 Introduction
We conduct a thorough experimental evaluation on multiple datasets that were not used during training, using powerful perceptual metrics such as LPIPS (Zhang et al., 2018) and DISTS (Ding et al., 2022) to assess image fidelity. SHARP improves image fidelity by substantial factors versus prior feedforward methods. In comparison to diffusion-based systems, SHARP delivers higher fidelity while reducing synthesis time by two to three orders of magnitude. Compared to the strongest prior method (Ren et al., 2025), SHARP reduces LPIPS by 25–34% and DISTS by 21–43% across the test datasets (in the zero-shot regime), while accelerating synthesis by three orders of magnitude and producing a 3D representation that supports high-resolution rendering of nearby views at 100 frames per second.
In summary, our contributions are as follows:
@@
Using our insights, we demonstrate that state-of-the-art high-resolution view synthesis is feasible in a purely regression-based framework.
@@ 2 Related Work
A number of works develop feedforward prediction models for view synthesis from a small number of nearby views. IBRNet (Wang et al., 2021) generalized image-based rendering across scenes using learned features and ray transformers. MVSNeRF (Chen et al., 2021) reconstructed neural radiance fields from a few input images via cost volume processing. LaRa (Chen et al., 2024) regressed an object-level radiance field from sparse input images. GS-LRM (Zhang et al., 2024) leveraged a transformer to predict a 3D Gaussian representation from posed sparse images. Our work focuses on view synthesis from a single image.
View synthesis from a single image. The introduction of larger datasets (Zhou et al., 2018; Tung et al., 2024) enabled the transition from scene-specific multi-view optimization to learning-based pipelines that can infer a plausible 3D representation from a single image. Zhou et al. (2016) synthesized novel views from a single image through appearance flow. Subsequent work developed variants of depth-based warping (Wiles et al., 2020; Jampani et al., 2021) and multiplane images (MPI) (Zhou et al., 2018; Tucker & Snavely, 2020). AdaMPI (Han et al., 2022) adapted multiplane images to diverse scene layouts through plane depth adjustment and depth-aware color prediction, trained using a warp-back strategy on single-view image collections. Khan et al. (2023) proposed Tiled Multiplane Images (TMPI), which splits an MPI into many small tiled regions with fewer depth planes per tile, reducing computational overhead while maintaining quality. Several recent methods have drawn inspiration from the success of transformers in modeling long-range dependencies, leveraging large transformer-based encoder-decoder architectures to infer scene structure and appearance implicitly for novel views (Hong et al., 2024; Jin et al., 2025).
@@
Diffusion models have emerged as powerful tools for novel view synthesis with sparse input, offering high-quality results through iterative denoising processes (Po et al., 2024). Watson et al. (2023) developed an early application to view synthesis. Zero-1-to-3 (Liu et al., 2023) demonstrated zero-shot view synthesis by fine-tuning diffusion models on 3D object datasets. iNVS (Kant et al., 2023) repurposed diffusion inpainters for view synthesis by combining monocular depth estimation with inpainting. NerfDiff (Gu et al., 2023) distilled a 3D-aware diffusion model into NeRF by synthesizing virtual views to improve rendering under occlusion. These early attempts focused on object-level view synthesis.
At the scene level, ViewCrafter (Yu et al., 2025b), ZeroNVS (Sargent et al., 2024), CAT3D (Gao et al., 2024), SplatDiff (Zhang et al., 2025), Stable Virtual Camera (Zhou et al., 2025), Bolt3D (Szymanowicz et al., 2025b), Wonderland (Liang et al., 2025), WonderWorld (Yu et al., 2025a), See3D (Ma et al., 2025) and Gen3C (Ren et al., 2025) all applied diffusion models to view synthesis from sparse image sets or a single image. The diffusion-based approach supports impressive image quality from faraway viewpoints, leveraging diffusion priors to synthesize plausible appearance even for views that have no overlap with the input. On the other hand, image quality from nearby views (corresponding to natural head motion or posture shifts) can be noticeably less sharp and photorealistic than the input, while the synthesis time can sometimes stretch into minutes. (Although Bolt3D (Szymanowicz et al., 2025b) makes impressive progress on the latter front.) In contrast, we aim for real-time rendering of maximally photorealistic high-resolution images from nearby views, supporting a headbox that allows for natural posture shifts while maintaining photographic quality. Our approach generates a high-resolution 3D representation that provides such experiences from single-image input in less than a second on a single GPU, supporting conversion of pre-existing photographs to photorealistic 3D during interactive browsing of a photo collection.
@@ 3.1 Overview
The input image $\mathbf{I}\in\mathbb{R}^{3\times H\times W}$ is fed into a pretrained Depth Pro image encoder $\varphi_{\text{enc}}$ to produce 4 intermediate feature maps $(\mathbf{f}_{i})_{i\in\{1,\dots,4\}}=\varphi_{\text{enc}}(\mathbf{I})$. As in Depth Pro (Bochkovskii et al., 2025), we resize the input image so that $H=W=1536$. We then feed the intermediate feature maps into the Depth Pro decoder $\varphi_{\text{dec}}$ to produce a monocular depth map $\mathbf{\hat{D}}=\varphi_{\text{dec}}((\mathbf{f}_{i})_{i\in\{1,\dots,4\}})$. Similar to Flynn et al. (2019), we duplicate the last convolutional layer of the decoder to produce a two-channel depth map $\mathbf{\hat{D}}\in\mathbb{R}^{2\times H\times W}$.
@@
While the initial Gaussians provide a reasonable starting point, they require substantial refinement to achieve high-fidelity rendering. The Gaussian decoder $\varphi_{\text{gauss}}$ takes as input the feature maps $(\mathbf{f}_{i})_{i\in\{1,\dots,4\}}$ and the input image $\mathbf{I}$, and outputs refinements $\Delta\mathbf{G}\in\mathbb{R}^{K\times 2\times H^{\prime}\times W^{\prime}}$} for all Gaussian attributes:
$\Delta\mathbf{G}=\varphi_{\text{gauss}}((\mathbf{f}_{i})_{i\in\{1,\dots,4\}},\mathbf{I}).$ (3.1)
@@
$\mathbf{G}_{\text{attr}}=\gamma_{\text{attr}}\bigg(\gamma_{\text{attr}}^{-1}(\mathbf{G}_{0,\text{attr}})+\eta_{\text{attr}}\Delta\mathbf{G}_{\text{attr}}\bigg).$ (3.2)
@@ 3.2 Network Architecture
Feature encoder. We base our feature encoder on the Depth Pro backbone (Bochkovskii et al., 2025). This encoder processes the input image $\mathbf{I}\in\mathbb{R}^{C\times H\times W}$} and produces four feature maps $(\mathbf{f}_{i})_{i\in\{1,\dots,4\}}$ at different resolutions. The Depth Pro backbone consists of two Vision Transformers (ViTs) (Dosovitskiy et al., 2021), one applied to a downscaled version of the input image and one applied to various image patches. The low-resolution image encoder and patch encoder each have 326M parameters. During training, we unfreeze the low-resolution image encoder to allow adaptation to the view synthesis task, while keeping the patch encoder and normalization layers frozen to preserve the pretrained feature extraction capabilities.
Depth decoder. Our depth decoder is based on the Dense Prediction Transformer (DPT) (Ranftl et al., 2021). We modify the original DPT decoder by duplicating the final convolutional layer to output two depth channels instead of one. Our decoder thus takes the feature maps $(\mathbf{f}_{i})_{i\in\{1,\dots,4\}}$ from the encoder and produces a two-layer depth map $\mathbf{\hat{D}}\in\mathbb{R}^{2\times H\times W}$. The first layer represents the primary visible surfaces, while the second layer may represent occluded regions and view-dependent effects. The depth decoder consists of multiple convolutional blocks with approximately 20M parameters. This module is fully unfrozen during training to optimize depth prediction for view synthesis.
Gaussian decoder. The Gaussian decoder predicts refinements for all Gaussian attributes. It has the same DPT architecture as the depth decoder but we replace the last upsampling block with a custom prediction head. The decoder takes as input the feature maps $(\mathbf{f}_{i})_{i\in\{1,\dots,4\}}$, the input image $\mathbf{I}$, and the predicted depth maps $\mathbf{\hat{D}}$. It outputs a tensor $\Delta\mathbf{G}\in\mathbb{R}^{K\times 2\times H^{\prime}\times W^{\prime}}$ that contains deltas for all Gaussian attributes: position (3 channels), scale (3 channels), rotation (4 channels), color (3 channels), and opacity (1 channel). This decoder has approximately 7.8M parameters and is trained from scratch. The high dimensionality of the output (approximately 16.5M values) enables fine-grained control over the Gaussian representation.
@@ 3.4 Training Objectives
where $\Omega$ denotes the set of all pixels $p$. We further use a perceptual loss (Johnson et al., 2016; Gatys et al., 2016; Suvorov et al., 2022) on novel views to encourage plausible inpainting:
@@
$\mathcal{L}_{\text{delta}}=\mathbb{E}_{i\sim\mathcal{I}}\left[\max\{\left|\Delta\mathbf{G}_{x}(i)\right|-\delta,0\}+\max\{\left|\Delta\mathbf{G}_{y}(i)\right|-\delta,0\}\right],$ (3.9)
@@
$\mathcal{L}_{\text{scale}}=\mathbb{E}_{p\sim\Omega}\left[|\mathbf{S}(p)-1|\right]\quad\text{ and }\quad\mathcal{L}_{\nabla\text{scale}}=\sum_{k=1}^{6}\mathbb{E}_{p\sim\Omega_{\downarrow k}}\left[|\nabla\mathbf{S}_{\downarrow k}(p)|\right].$ (3.11)
@@ 4 Experiments
We employ LPIPS (Zhang et al., 2018) and DISTS (Ding et al., 2022) to quantitatively assess the quality of novel view synthesis. We focus primarily on these perceptual metrics, since older pointwise metrics such as PSNR and SSIM can be overly sensitive to small translations, where even a 1% shift can lead to catastrophic drops in scores despite visually similar results. (See the supplement for an illustration. We also list PSNR and SSIM numbers in the supplement for completeness.) Since we are interested in sharp high-resolution view synthesis, we evaluate all methods on the full-resolution ground truth. If a method generates results at lower resolution, we resize the output image to the input resolution before evaluation. If a method crops the input and generates cropped results (Ren et al., 2025; Yu et al., 2025b; Jin et al., 2025; Zhou et al., 2025), we evaluate against correspondingly cropped ground-truth images.
@@
We compare SHARP to the following state-of-the-art methods: Flash3D (Szymanowicz et al., 2025a), which is based on 3D Gaussians; TMPI (Khan et al., 2023), which uses multi-plane images; LVSM (Jin et al., 2025), which is based on image-to-image regression; and Stable Virtual Camera (SVC) (Zhou et al., 2025), ViewCrafter (Yu et al., 2025b), and Gen3C (Ren et al., 2025), which employ diffusion models.
@@ 5 Conclusion
One clear opportunity for future work is to extend the methodology to support photorealistic synthesis of faraway views without compromising the fidelity of nearby views or the benefits of fast interactive synthesis. This may call for judicious integration of diffusion models (Po et al., 2024), possibly with the aid of distillation for reducing synthesis latency (Yin et al., 2024). With diffusion models, a unified view synthesis routine for single-view, multi-view, and video input (Ren et al., 2025; Ma et al., 2025) may emerge as a versatile generalization. Another interesting avenue is a principled treatment of view-dependent and volumetric effects (Verbin et al., 2024).
@@ A.1 Attribute Specific Activation
Activation functions $\gamma$ and their corresponding scale factors $\eta$ in Eq. 3.2 are specified below:
@@ A.3 View Frustum Masking
We implement a view frustum masking technique to address ambiguity in view synthesis, since regions occluded in the original view have multiple plausible reconstructions. By using depth information to determine which regions in the new view correspond to points visible in the original view, we apply supervision only where ground truth is reliable.
@@
A.4 The Perceptual Loss
Here we detail the challenges and our solutions in incorporating the perceptual loss. We employ the perceptual loss aimed at improving inpainting (Suvorov et al., 2022). Similar to its application in the image domain, we initially applied the loss only in the occluded image patches of the novel view; however, through experiments, we observed that applying the loss to the entire rendered image resulted in more plausible details and fewer artifacts in general, even in the non-occluded foreground regions, as seen in Figure 9.
@@ B.1 Synthetic data
To further enhance scene diversity and complexity, the framework supports random placement of various object types, including thin structures, transparent materials, and reflective surfaces, across a wide range of spatial configurations. It also offers fine-grained control over camera parameters such as position, orientation, and focal length, as well as detailed illumination settings. Lighting setups include physically-based direct light sources, with variations in direction, intensity, and color temperature. We also use high-dynamic-range (HDR) environment maps, which are sampled from a curated collection of high-resolution HDRIs. This combination enables realistic global illumination effects under diverse and physically plausible lighting conditions.
For each scene, we identify one object of interest and position a ring containing 10 virtual cameras around it at varying distances and angles. The cameras are arranged in concentric circles such that the cameras are no more than 60 cm apart, simulating a multi-view capture setup. This allows the dataset to capture the same object or scene element from diverse perspectives. All images are rendered using the V-Ray physically based rendering engine, ensuring photorealistic lighting and material interactions. The final dataset consists of approximately 700K unique rendered scene instances, each with 11 rendered views, totaling around 8M images at $1536\times 1536$ or $2048\times 2048$ resolutions.